# A Reanalysis of the FDA’s Benefit-Risk Assessment of Moderna’s mRNA-1273 COVID Vaccine: For 18–25-Year-Old Males, Risks Exceeded Benefits Relative to Hospitalizations

### v1

@@
Abstract: The Food and Drug Administration (FDA) conducted a benefit-risk assessment for Moderna’s COVID vaccine mRNA-1273 prior to its full approval (announced 1/31/2022). The FDA focused its assessment on males of ages 18–64 years because the agency’s risk analysis was limited to vaccine-attributable myocarditis/pericarditis (VAM/P) given the excess risk among males. The FDA analysis concluded that vaccine benefits clearly outweighed risks, even for 18–25-year-old males (those at highest VAM/P risk). In this paper, we reanalyze the FDA benefit-risk assessment using information available through the third week of January 2022 and focusing on 18–25-year-old males. We use the FDA’s framework but extend its model by accounting for protection derived from prior COVID infection, finer age-stratification in COVID-hospitalization rates, and incidental hospitalizations (those of patients who test positive for COVID but are being treated for something else). We also use more realistic projections of Omicron-infection rates and more accurate rates of VAM/P. With hospitalizations as the principal endpoint of the analysis (those prevented by vaccination vs. those caused by VAM/P), we find a net harm: vaccine risks outweighed benefits for 18–25-year-old males, except in scenarios projecting implausibly high Omicron-infection prevalence. Comparing the projections of FDA’s “most likely scenario” to ours, we find that the FDA’s benefit-risk ratio, 43.33, is over 60 times more favorable than ours, 0.67. Our assessment suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myocarditis/pericarditis alone compared to COVID hospitalizations prevented (over a five-month period of vaccine protection assumed by the FDA). The preceding assessment is based on assumptions drawn from data available before 1/22/22. Moreover, these assumptions are validated by subsequently available data.
@@ 1. Introduction
The Food and Drug Administration (FDA) issued emergency-use authorizations for mRNA COVID-19 vaccines from Pfizer (BNT162b2) and Moderna (mRNA-1273) in December 2020 [2]. Once the mRNA COVID-19 vaccines were given to large populations, it became evident that the vaccines could cause myocarditis and pericarditis (hereafter myo/pericarditis), inflammation of the heart and of the membrane surrounding the heart, respectively, with studies indicating the highest risk among males younger than 40 years ([3, 4, 5, 6, 7]).
@@
The FDA has published two benefit-risk analyses of primary series mRNA vaccination, the first for BNT162b2 [8] and the second for mRNA-1273 [9]. The FDA presented its analyses through a number of scenarios involving different modeling assumptions to account for uncertainties associated with pandemic dynamics as well as with the benefits and risks of vaccination. To assess vaccine benefits, the FDA estimated vaccine-preventable COVID-19-cases, hospitalizations, intensive care unit (ICU) admissions, and deaths [8, 9]. To assess vaccine risks, the FDA estimated vaccine-attributable myo/pericarditis (VAM/P) cases, hospitalizations, ICU admissions, and deaths [8, 9]. The FDA concluded that the benefits of the mRNA vaccines outweighed the risks of VAM/P even for males in its highest-risk age-groups: 16–17 years for BNT162b2 [8] and 18–25 years for mRNA-1273 [9].
@@
However, both of the FDA’s benefit-risk assessments omitted important inputs in the analysis of the hospitalizations-prevented benefit of vaccination as well as in the analysis of VAM/P risk. We focus mainly on the FDA’s benefit-risk analysis for Moderna’s mRNA-1273 to simplify our exposition—there are significant differences in modeling assumptions in the FDA’s Pfizer and Moderna assessments. Also, VAM/P risk is substantially greater for mRNA-1273 than for BNT162b2 [3, 5, 6, 10].
@@
For the FDA’s assessment of mRNA-1273, benefits of vaccination were overestimated. First, the FDA did not account for the protection conferred by prior infection; and, at the time of the FDA’s Moderna analysis (mid-January 2022), it is likely that over 70% of 18–29 year-olds had been infected by COVID-19—see Supplement 1 for details. Second, the FDA did not account for incidental COVID-19 hospitalizations. Third, the FDA overestimated benefits for young males by assuming the COVID-hospitalization rate for 18–25 year-olds is equivalent to that for 26–35 year-olds and 36–45 year-olds. Finally, the FDA assumed in its most likely scenario that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months.
By contrast, the FDA underestimated VAM/P risks for mRNA-1273. The agency omitted dose-one mRNA-1273 VAM/P events [9, Section 2.3.1] (but included dose-one events for its BNT162b2 assessment [8, Section 2.3.1]). Moreover, the FDA’s analysis is inconsistent with evidence from the CDC that VAM/P risk for mRNA-1273 was over twice that for BNT162b2 [10, p. 25]. Finally, the FDA relied on internal FDA databases that yielded mRNA-1273 VAM/P rates substantially lower than those found in major international studies [3, 5, 6] as well as a U.S. study [4].
@@
In this study, our primary aim is to extend and improve the FDA’s benefit-risk model for mRNA-1273 vaccination by accounting for prior-infection protection, incidental COVID-19 hospitalizations, and age-specific infection-hospitalization rates, as well as for a more accurate assessment of VAM/P risk. We apply our model using data available to the FDA at the time of its benefit-risk analysis for mRNA-1273, which we take to be the 3rd week of January 2022 for reasons described Section 2.1 below. We also analyze benefits and risks of vaccination for specific subpopulations such as males 18–25 having prior-infection protection and no comorbidities. Our modeling not only addresses the shortcomings in the FDA’s assessment of the mRNA-1273 vaccine but also presents a framework for future benefit-risk assessments of vaccination.
@@ 2. The FDA’s mRNA-1273 Benefit-Risk Assessment
In [9], the FDA describes its benefit-risk modeling of Moderna’s mRNA-1273 (Spikevax) for males ages 18 to 64 years. Outcomes are provided via six scenarios reflecting uncertainties in three major model inputs: COVID-19 incidence, vaccine effectiveness against cases (VE) as well as hospitalization (VEH), and vaccine-attributable myo/pericarditis (VAM/P) rate. At the time of the FDA’s mRNA-1273 analysis, the Centers for Disease Control and Prevention (CDC) was recommending a 5-month interval between primary-series completion and a booster dose. Consequently, the FDA evaluated benefits and risks over a 5-month period, which we take to be 1/1/22 – 5/31/22. The FDA assumed that all vaccine effectiveness estimates remained constant over this 5-month evaluation period. The FDA’s major modeling assumptions are summarized in Table 1 below.
Table 1. The FDA’s mRNA-1273 model-input assumptions for 18–25-year-old males [9, Table 1].
Scenario COVID-19 Incidence Vaccine Effectiveness VAM/P Rate Assumptions Per Million 2nd Vaccinations*
$1$ Highest 2021 Incidence, 12/25–12/31 Omicron Dominant; VE = 30%, VEH=72% 128 cases, 110 hospitalizations
@@ 6 Same as Scenario 1 Same as Scenario 1 241 cases, 207 hospitalizations
*From [9, Table 2].
Using the modeling assumptions described in the first row of the Table 1, with incidence data drawn from [9, Table 3], the FDA concluded for Scenario 1, its “most likely scenario,” that vaccinating one million males ages 18–25 years with two doses would prevent, over 5 months, 82,484 COVID-19 cases, 4,766 COVID-19 hospitalizations, 1,144 ICU admissions, and 51 deaths due to COVID-19, while risking 128 VAM/P cases, 110 VAM/P hospitalizations, 0 ICU admissions, and 0 deaths due to VAM/P [9, Table 2]. However, the FDA’s model overestimates benefits and underestimates risks.
@@ 2.1. The Time of the FDA’s mRNA-1273 Analysis
Although the FDA does not define in [9] the time of its benefit-risk analysis for mRNA-1273, a natural choice is the 3rd week of January 2022, which concluded 10 days prior to the FDA’s announcement of the full approval of mRNA-1273 on 1/31/2022 [12]. This choice is natural because it is consistent with the FDA’s time-of-analysis definition for its BNT162b2 assessment: the 2nd week of August 2021 [8, Section 2.1], which concluded nine days prior to the FDA’s announcement of the full approval of BNT162b2 on 8/23/2021 [11]. Consequently, our reanalysis of the FDA’s benefit-risk assessment [9] for the general population of 18–25-year-old males is based on assumptions derived from information available before 1/22/2022. Subsequent information validating our assumptions will often be provided but only via footnotes. Starting in Section 4.2, we no longer restrict ourselves to basing modeling assumptions only on information available before 1/22/2022; in this section, our intent is to demonstrate how our model might be modified to incorporate a comorbidity (obesity) as a risk modifier.
Starting in Section 4.2, where we explore how vaccination’s benefit-risk picture changes with respect to Body Mass Index (BMI), we no longer restrict ourselves to basing modeling assumptions only on information available before 1/22/2022.
@@ 2.2. The FDA’s mRNA-1273 Benefits Assessment
[…]
@@
For Scenario 1, its most likely scenario, and two others, the FDA assumed that peak 2021 Omicron-case rates (late December 2021) would be sustained for 5 months, resulting in an exceptionally high projection of total COVID-19 cases among the unvaccinated; for example, $C=274,947$ for 18–25-year-old males. Similarly, for its most likely scenario (and two others), the FDA assumed that the CDC’s reported late-December 2021 Omicron hospitalization rate for unvaccinated 18–45-year-old males would be sustained for 5 months at approximately 43.54 cases per million daily [9, Table 3] and assumed that the same rate would apply to the younger 18–25 bracket, arriving at $H=6,619$ hospitalizations. Thus, for its most likely scenario (and two others), the FDA predicted that full mRNA-1273 vaccination of 1 million 18–25-year-old males would prevent $274,947\cdot 0.30\approx 82,484$ COVID cases and $6619\cdot 0.72\approx 4,766$ COVID hospitalizations [9, Table 2].
To obtain its projections of ICU admissions prevented by vaccination, the FDA used CDC data to estimate age-based percentages of males hospitalized with COVID going to ICU [9, Table 3]. According to Table 3 of [9], 24% of 18–25-year-old males admitted with COVID require ICU care; thus, the FDA predicted that, under its most likely scenario, fully vaccinating 1 million males in this age group with mRNA-1273 would prevent approximately 1,144 ICU admissions (24% of 4,766)[9, Table 2]. The FDA’s projections of deaths prevented by vaccination utilized the same assumptions as hospitalizations prevented. Specifically, for its most likely scenario, the FDA assumed the CDC’s reported late-December 2021 Omicron death rate for unvaccinated males 18–25 would be sustained for 5 months resulting in a total of about 71 deaths per million and assumed vaccine effectiveness against death to be the same as VEH, estimating $71\cdot 0.72\approx 51$ COVID deaths prevented per million vaccinations.
@@
$\bullet$ The FDA did not account for the protection conferred by prior infection. The FDA’s not accounting for prior-infection protection in its mRNA-1273 assessment is difficult to justify. A study of SARS-CoV-2–specific immune memory found “Substantial immune memory is generated after COVID-19,” with about 95% of subjects retaining memory at  6 months post infection [15]. A meta-analysis based on studies published December 2020 through August 2021 demonstrated that protection from prior COVID infection “is, at least, equivalent to the protection afforded by complete vaccination of COVID-naïve populations” [16]. The CDC acknowledged in a science brief released 10/29/21 that “The immunity provided by vaccine and prior infection are both high” [17], while another CDC study found that during the Delta wave unvaccinated persons with prior COVID infection had significantly greater protection against infection and hospitalization than fully vaccinated, COVID-naïve persons [18]. Moreover, by the end of December 2021, the FDA acknowledged, “The omicron variant has significantly more mutations than previous SARS-CoV-2 variants, particularly in its S-gene, the gene that encodes the virus’s spike protein”[19], suggesting that prior infection, which entails exposure to all SARS-CoV-2 proteins (such as the nucleoprotein), would likely yield greater protection against Omicron than would mRNA-1273 vaccination, which entails exposure only to the spike protein of the ancestral strain of SARS-CoV-2. The summary of findings of a study published 7/20/21 concludes, “CD4+ T cell responses equally target several SARS-CoV-2 proteins, whereas the CD8+ T cell responses preferentially target the nucleoprotein, highlighting the potential importance of including the nucleoprotein in future vaccines. Taken together, these results suggest that broad and effective immunity may persist long-term in recovered COVID-19 patients[20]. [^1] Consistent with the immunological considerations discussed in this paragraph, a study from Qatar found prior infection provides 52.2% effective protection against symptomatic Omicron infection and 91% effective protection against “severe, critical, or fatal” Omicron infection; whereas, the study’s corresponding effectiveness estimates for 2 doses of mRNA-1273 are 2.2% and 66.3%, respectively [21, Figure 2C, 2D] (study period 12/23/21 – 2/21/22). In Section 7 of Supplement 1, we explain how the meta-analyses and systematic reviews [22, 23, 24, 25] provide convincing evidence that prior infection provided greater protection against Omicron during the evaluation period than did two-dose vaccination among the COVID-naïve.
@@
Additional reasons why the FDA’s not accounting for prior-infection protection is difficult to justify include (i) according to the CDC, 54.9% of those 18–49 years old had been infected by COVID-19 by the end of September 2021 [26]; (ii) a South African study [27] released 12/14/21 suggested overall effectiveness of prior-infection protection against Omicron to be about 47% (see Supplement 1), while data from England [28] released 12/22/21 suggests that prior infection among the unvaccinated provides significant protection from Omicron hospitalization (see Supplement 1); and (iii) the FDA adopted a scenario-based structure for its benefit-risk assessment to account for “the impact of uncertainty associated with pandemic dynamics” [9, Abstract], yet the agency did not consider prior-infection protection in any of its six scenarios.
@@
$\bullet$ The FDA did not account for incidental COVID-19 hospitalizations. At the time of its mRNA-1273 assessment, it was well known that incidental-hospitalization rates were substantial. The day after CDC director R. Walensky [29] stated during a 1/9/22 interview “[U]p to 40 percent of the patients who are coming in with COVID-19 are coming in …with something else,” [^2] We searched unsuccessfully for published CDC data on incidental COVID-hospitalization rates. Figure 2 of a large U.S. study suggests an incidental rate for 0–18 year-olds typically well above 50% during the study period 3/1/2020–1/31/22 and for those 19–64 an incidental rate above 40% for 1/1/2022–1/31/2022 [30]. Canada’s province of Ontario started publishing incidental rates, with rates for January 8–15 exceeding 45% [31]. There is reason to believe incidental rates are especially high for the young—see Supplement 1.
$\bullet$ The FDA overestimated the proportion of projected COVID cases requiring hospitalization. One way to gauge the degree of implausibility of the FDA’s evaluation-period hospitalization projections is to consider the corresponding case-hospitalization rates:
@@
Table 2. Case-hospitalization-rate projections for unvaccinated 18–25-year-old males over the evaluation period 1/1/22 – 5/31/22 for the Omicron-based scenarios of [9, Table 2].*
@@
$5$ 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
$6$ 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
@@
The FDA’s projected case-hospitalization rates are all over four times the Omicron case-hospitalization rate ($0.58\%\approx 56/9,585$) for the unvaccinated of all ages based on Omicron cases in England tested between 1 December 2021 and 14 December 2021, inclusive [28, Table 3]; moreover, all are well over two times the case-hospitalization rate (0.92%) for the unvaccinated of ages 18–29 in Ontario over the period 12/14/20 – 12/12/21 [32, Table A1a]. [^3] The FDA case-hospitalization rates in Table 2 are all well over 5 times the Omicron rate 0.42% for the unvaccinated of all ages reported in Table S10 of [33] and based on data from the period 12/15/21 – 1/17/22. Based on information available at the time of the FDA’s mRNA-1273 assessment, one would expect the COVID-case hospitalization rate for the period 12/14/20 – 12/12/21 would be higher than that for Omicron; e.g., according to [34, p. 4], “The risk of hospital admission from emergency departments with Omicron was approximately one-third of that for Delta (Hazard Ratio 0.33, 95% CI: 0.30 to 0.37)”; see Supplement 1 for more examples. Furthermore, we expect the hospitalization rate for the age-group 18–25 to be lower than that for 18–29 [35]. To some extent, FDA’s exaggerated case-hospitalization rates for the 18–25 age range result from the agency’s assumption that hospital-admission rates were uniform for ages 18–45, contradicting the CDC’s own models [36]. Also, to some extent, the FDA’s exaggerated case-hospitalization rates result from the CDC’s inclusion of incidental COVID hospitalizations in its accounting. Our source for the 0.58% rate mitigated the incidental-hospitalization issue [28, p. 10], while our source [32] for the 0.92% rate did not.

### v3

# A reanalysis of the FDA’s benefit-risk assessment of Moderna’s mRNA-1273 COVID vaccine stratified not only based on age and sex but also on prior-infection and comorbidity status

@@
Abstract. We demonstrate that the outcome of a benefit-risk assessment for vaccination may be dramatically impacted by accounting for the benefits derived from prior infection by the disease against which the vaccine is targeted. The United States Food and Drug Administration (FDA) conducted a benefit-risk assessment for Moderna’s COVID vaccine mRNA-1273 prior to its full approval, announced 1/31/2022. The FDA’s assessment focused on males of ages 18–64 years because the agency’s risk analysis was limited to vaccine-attributable myocarditis/pericarditis (VAM/P) given the excess risk among males. The FDA’s analysis concluded that vaccine benefits clearly outweighed risks, even for 18–25-year-old males (those at highest VAM/P risk). We reanalyze the FDA’s benefit-risk assessment using information available through the third week of January 2022 and focusing on 18–25-year-old males. We use the FDA’s framework but extend its model by accounting for protection derived from prior COVID infection, finer age-stratification in COVID-hospitalization rates, and incidental hospitalizations (those of patients who test positive for COVID but are being treated for something else). We also use more realistic projections of Omicron-infection rates and more accurate rates of VAM/P. With hospitalizations as the principal endpoint of the analysis (those prevented by vaccination vs. those caused by VAM/P), our model finds vaccine risks outweighed benefits for 18–25-year-old males, except in scenarios projecting implausibly high Omicron-infection prevalence. Our assessment suggests that mRNA-1273 vaccination of 18–25-year-old males generated between 16% and 63% more hospitalizations from vaccine-attributable myocarditis/pericarditis alone compared to COVID hospitalizations prevented (over a five-month period of vaccine protection assumed by the FDA). The preceding assessment derives from model inputs based on data available at the time of the FDA’s mRNA-1273 assessment. Moreover, these inputs as well as model outputs are validated by subsequently available data.
@@ 1. Introduction
The Food and Drug Administration (FDA) issued emergency-use authorizations for mRNA COVID-19 vaccines from Pfizer (BNT162b2) and Moderna (mRNA-1273) in December 2020 [2]. Once the mRNA COVID-19 vaccines were given to large populations, it became evident that they could cause myocarditis or pericarditis (hereafter myo/pericarditis), inflammation of the heart or of the membrane surrounding the heart, respectively, with studies indicating the highest risk among males younger than 40 years [3, 4, 5, 6, 7, 8, 9].
@@
The FDA has published two benefit-risk analyses of primary series mRNA vaccination, the first for BNT162b2 [10] and the second for mRNA-1273 [11]. The FDA presented its analyses through scenarios involving different modeling assumptions to account for uncertainties in pandemic dynamics as well as in the benefits and risks of vaccination. To assess vaccine benefits, the FDA estimated vaccine-preventable COVID-19 cases, hospitalizations, intensive-care unit (ICU) admissions, and deaths [10, 11]. To assess vaccine risks, the FDA estimated vaccine-attributable myo/pericarditis (VAM/P) cases, hospitalizations, ICU admissions, and deaths [10, 11]. The FDA concluded that the benefits of the mRNA vaccines outweighed VAM/P risk even for males in its highest-risk age-groups: 16–17 years for BNT162b2 [10] and 18–25 years for mRNA-1273 [11].
@@
However, these FDA benefit-risk assessments omitted important inputs in the analysis of both the hospitalizations-prevented benefit of vaccination and VAM/P risk. We focus mainly on the FDA’s analysis for Moderna’s mRNA-1273 to simplify our exposition—there are significant differences in modeling assumptions in the FDA’s Pfizer and Moderna assessments. Also, VAM/P risk is substantially greater for mRNA-1273 than for BNT162b2 [3, 4, 5, 6, 7, 8, 9].
@@
For the FDA’s assessment of mRNA-1273, benefits of vaccination were overestimated. First, the FDA did not account for the protection conferred by prior infection; and, at the time of the FDA’s Moderna analysis (mid-January 2022), likely over 70% of 18–25 year-olds in the U.S. had been infected by COVID-19—see Supplement 1. Second, the FDA did not account for incidental COVID-19 hospitalizations. Third, the FDA overestimated benefits for young males by assuming the COVID-hospitalization rate for 18–25 year-olds is equivalent to that for 26–35 year-olds and 36–45 year-olds. Finally, the FDA assumed in its most likely scenario that case and hospitalization rates would remain constant at peak 2021 Omicron levels for 5 months.
By contrast, the FDA underestimated mRNA-1273 VAM/P risks; see, e.g.,Table 3 of Section 2.3.
Here, our primary aim is to demonstrate how to incorporate the protection derived from prior infection into a population-level quantitative benefit-risk model for a two-dose vaccination series. Our model has hospitalizations as its endpoint, comparing the hospitalizations-prevented benefit of vaccination to the hospitalization risk attributable to vaccination. We illustrate the utility of our model by applying it to extend and improve the FDA’s mRNA-1273 assessment relative to its hospitalizations endpoint, which can be viewed as the most important of the agency’s endpoints owing to the public-health concern that during waves of the COVID pandemic hospitals might be so overwhelmed that they would not be able to deliver high quality care to their patients.
@@
In addition to prior-infection protection, our modeling accounts for incidental hospitalizations, age-specific infection-hospitalization rates, and a more accurate assessment of VAM/P risk. We apply our model using data available to the FDA at the time of its benefit-risk analysis for mRNA-1273, assumed to be the 3rd week of January 2022 for reasons described in Section 2.1. We present evidence, using real-world data, that our model’s estimates of the hospitalizations-prevented benefit of mRNA-1273 vaccination are significantly more accurate than those the FDA obtained in [11]. We also analyze benefits and risks of vaccination for specific subpopulations such as males 18–25 having prior-infection protection and no comorbidities. Our model can be adapted to produce for any vaccine designed for deployment on a population-wide scale a benefit-risk assessment stratified not only based on age and sex but also on prior-infection and comorbidity status.
@@ 2. The FDA’s mRNA-1273 Benefit-Risk Assessment
In [11], the FDA describes its benefit-risk modeling of Moderna’s mRNA-1273 for males ages 18 to 64 years. Outcomes are provided via six scenarios reflecting uncertainties in three major model inputs: COVID-19 incidence, vaccine effectiveness against cases (VE) as well as hospitalization (VEH), and vaccine-attributable myo/pericarditis (VAM/P) rate. At the time of the FDA’s mRNA-1273 analysis, the CDC was recommending a 5-month interval between primary-series completion and a booster dose. Consequently, the FDA evaluated benefits and risks over a 5-month period, which we take to be 1/1/22 – 5/31/22. The FDA assumed that all vaccine effectiveness estimates remained constant over this 5-month evaluation period. The FDA’s major modeling assumptions are summarized in Table 1 below.
Table 1. The FDA’s mRNA-1273 model-input assumptions for 18–25-year-old males [11, Table 1].
Scenario COVID-19 Incidence Vaccine Effectiveness VAM/P Rate Per Million 2nd doses [11, Table 2]
11 Highest 2021 Incidence, 12/25–12/31 Omicron Dominant; VE = 30%, VEH=72% 128 cases, 110 hospitalizations
@@ 6 Same as Scenario 1 Same as Scenario 1 241 cases, 207 hospitalizations
Using inputs from first row of Table 1 and incidence data from [11, Table 3], the FDA concluded for Scenario 1, its “most likely scenario,” that vaccinating one million males ages 18–25 years with two doses would prevent, over 5 months, 82,484 COVID-19 cases, 4,766 COVID-19 hospitalizations, 1,144 ICU admissions, and 51 deaths due to COVID-19, while risking 128 VAM/P cases, 110 VAM/P hospitalizations, 0 ICU admissions and 0 deaths from VAM/P [11, Table 2]. .
@@ 2.1. The Time of the FDA’s mRNA-1273 Analysis
Although the FDA does not define in [11] the time of its benefit-risk analysis for mRNA-1273, a natural choice is the 3rd week of January 2022, which concluded 10 days before the FDA’s announcement of the full approval of mRNA-1273 on 1/31/2022 [12]. This choice is consistent with the FDA’s time-of-analysis definition for its BNT162b2 assessment: the 2nd week of August 2021 [10, Section 2.1], which concluded nine days before the FDA’s announcement of the full approval of BNT162b2 on 8/23/2021 [13]. Consequently, our reanalysis of the FDA’s benefit-risk assessment [11] is based on model inputs derived from information available before 1/22/2022. Subsequent information validating our input choices will often be provided but only via footnotes. Starting in Section 4.2, we discontinue basing our modeling on information only available before 1/22/2022; in Section 4.2, our intent is to show how our model can be modified to incorporate a comorbidity (obesity) as a risk modifier.
@@ 2.2. The FDA’s mRNA-1273 Benefits Assessment
$C\cdot VE\ \text{cases prevented and}\ H\cdot VEH\ \text{hospitalizations prevented},$
@@
For Scenario 1, its most likely scenario, and two others, the FDA assumed that peak 2021 Omicron-case rates (late December 2021) would be sustained for 5 months, resulting in an exceptionally high projection of COVID-19 cases among the unvaccinated; e.g., $C=274,947$ for 18–25-year-old males. Similarly, for these scenarios, the FDA assumed that the CDC’s reported late-December 2021 Omicron hospitalization rate for unvaccinated 18–45-year-old males would be sustained for 5 months at approximately 43.54 cases per million daily [11, Table 3] and assumed that the same rate would apply to the younger 18–25 bracket, arriving at $H=6,619$ hospitalizations. Thus, for these scenarios, the FDA predicted that full mRNA-1273 vaccination of 1 million 18–25-year-old males would prevent $274,947\cdot 0.30\approx 82,484$ COVID cases and $6619\cdot 0.72\approx 4,766$ COVID hospitalizations over the 5-month evaluation period [11, Table 2].
@@
The FDA did not account for the protection conferred by prior infection. The FDA’s disregarding prior-infection protection in its mRNA-1273 assessment is difficult to justify. Dan et al. found “About 95% of subjects retained immune memory at $\sim$6 months after [COVID] infection” [14]. A meta-analysis based on studies published December 2020 through August 2021 demonstrated that protection from prior COVID infection “is, at least, equivalent to the protection afforded by complete vaccination of COVID-naïve populations” [15]. The CDC acknowledged in a science brief released 10/29/21 that “The immunity provided by vaccine and prior infection are both high” [16], while another CDC study found that during the Delta wave unvaccinated persons with prior COVID infection had significantly greater protection against infection and hospitalization than did fully vaccinated, COVID-naïve persons [17]. Moreover, in December 2021, the FDA acknowledged, “The omicron variant has significantly more mutations than previous SARS-CoV-2 variants, particularly in its S-gene, the gene that encodes the virus’s spike protein”[18], suggesting that prior infection, entailing exposure to all SARS-CoV-2 proteins (such as the nucleoprotein), would likely yield greater protection against Omicron than would mRNA-1273 vaccination, which entails exposure only to the spike protein of the ancestral strain of SARS-CoV-2. Among the findings of a study of SARS-CoV-2 specific B and T cell memory is “CD8+ T cell responses preferentially target the nucleoprotein, highlighting the potential importance of including the nucleoprotein in future vaccines” [19]. [^1] Consistent with the evidence presented in this paragraph, a Qatari study found prior infection provided 52.2% effective protection against symptomatic Omicron infection and 91% effective protection against “severe, critical, or fatal” Omicron, with corresponding effectiveness estimates for 2 doses of mRNA-1273 being 2.2% and 66.3%, respectively [20, Figure 2C, 2D] (study period 12/23/21 – 2/21/22). Findings from the meta-analyses and systematic reviews [21, 22, 23, 24] provide convincing evidence that prior infection provided greater protection against Omicron during the evaluation period than did two-dose vaccination among the COVID-naïve [Supplement 1, Section 7].
@@
Further evidence of the need to account for prior-infection protection includes (i) a CDC estimate that 54.9% of 18–49 year-olds had been infected by COVID-19 by 9/30/2021 [25], (ii) a South African study [26] released 12/14/21 suggesting overall effectiveness of prior-infection protection against Omicron to be about 47% [Supplement 1, Section 4], and (iii) data from England [27] released 12/22/21 suggested that prior infection among the unvaccinated provides significant protection from Omicron hospitalization [Supplement 1, Section 5].
@@
$\bullet$ The FDA did not account for incidental COVID-19 hospitalizations. At the time of its mRNA-1273 assessment, incidental-hospitalization rates were well known to be substantial. On 1/9/22, the director of the CDC stated, “[U]p to 40 percent of the patients who are coming in with COVID-19 are coming in …with something else” [28]. [^2] We searched unsuccessfully for published CDC incidental-COVID-hospitalizations data. Figure 2 of a large U.S. study suggests an incidental rate for 0–18 year-olds typically well above 50% during its study period 3/1/2020–1/31/22 and for those 19–64 an incidental rate above 40% for 1/1/2022–1/31/2022 [29]. On 1/10/22, Ontario began publishing incidental rates, with rates for January 8–15 exceeding 45% [30]. There is evidence that incidental rates are especially high for the young. For example, a major study [31] found, “As of May 2021, the cumulative rate for hospitalization associated with SARS-CoV-2 infection was 35.9 per 10,000 children” and “When limiting the analysis to include only patients with COVID-19 who required therapeutic interventions, the hospitalization rate decreased 5.5-fold to 6.5 per 10,000 children,” which suggests an incidental rate of over 80% “for children” (persons 0-17-years-old). See Supplement 1 for additional incidental-hospitalization data.
$\bullet$ The FDA overestimated the proportion of projected COVID cases requiring hospitalization. All of the FDA’s projected case-hospitalization rates, which are described in Table 2 below, are over four times the Omicron case-hospitalization rate ($0.58\%\approx 56/9,585$) for the unvaccinated of all ages based on Omicron cases in England tested between 1 December 2021 and 14 December 202 [27, Table 3]; moreover, all are well over twice the case-hospitalization rate (0.92%) for the unvaccinated of ages 18–29 in Ontario 12/14/20 – 12/12/21 [32, Table A1a]. [^3] The FDA case-hospitalization rates in Table 2 are all well over 5 times the Omicron rate 0.42% for the unvaccinated of all ages reported in Table S10 of [33] (data from 12/15/21 – 1/17/22). Based on information available at the time of the FDA’s mRNA-1273 assessment, we expect the COVID-case hospitalization rate for the period 12/14/20 – 12/12/21 to be higher than that for Omicron; e.g., “The risk of hospital admission from emergency departments with Omicron was approximately one-third of that for Delta (Hazard Ratio 0.33, 95% CI: 0.30 to 0.37)” [34, p. 4]; see Supplement 1 for more examples. Furthermore, we expect the hospitalization rate for those 18–25 to be lower than that for 18–29 [35]. Factors contributing to the FDA’s overestimating case-hospitalization rates for the 18–25 age range include (i) assuming hospital-admission rates were uniform for ages 18–45, contradicting CDC models [36], and (ii) including incidental COVID hospitalizations. Our source for the 0.58% rate mitigated the incidental-hospitalization issue [27, p. 10], while our source [32] for the 0.92% rate did not.
@@
Table 2. The FDA’s case-hospitalization-rate projections for unvaccinated 18–25-year-old males over the evaluation period 1/1/22 – 5/31/22 for the Omicron-based scenarios of [11, Table 2].*
@@
55 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$
66 274,947 6,619 $\frac{6619}{274,947}\approx 2.41\%$